OpenStudy (anonymous):

The equation of the curve whose slope at point (x,y) is x^2-2 and which contains the point (-1,3) is...

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@thomaster

OpenStudy (amistre64):

what is our 'antiderivative'? or just integrate

OpenStudy (anonymous):

1/3x^3-2x

OpenStudy (amistre64):

+C we need to determine C that satisfies the initial condition

OpenStudy (anonymous):

oh yeah

OpenStudy (amistre64):

so y = 1/3 x^3 - 2x + C, when x=-1, y=3 what should C be?

OpenStudy (anonymous):

-4/3?

OpenStudy (amistre64):

maybe :) 3 = -1/3 +2 + C 1 = -1/3 + C i dont think its -4/3

OpenStudy (anonymous):

x is actually 1, y is actually -3 I typed it wrong sorry

OpenStudy (amistre64):

pfft, then your on your own :)

OpenStudy (anonymous):

y=1/3x^3-2x-4/3, thanks for the help :)

OpenStudy (amistre64):

yep ;)